# Quasinormal modes of rotating black holes in shift-symmetric Einstein-scalar-Gauss-Bonnet theory

### v1

@@ 1 Introduction
There are various reasons to expect that GR will be superseded by a more complete theory of gravity. Thus in recent years numerous alternative theories of gravity have been suggested, driven largely by cosmological considerations [6]. Of course, any viable gravitational theory should conform with all experiments and observations [7]. But currently the strong gravity sector is still in need of far more testing [8]. Here the next generation of gravitational wave detectors should play an important role.
After the merger of two black holes the newly formed highly excited black hole undergoes ringdown where it emits damped sinusoidal gravitational waves, while it evolves towards a stationary state. These waves are associated with the quasinormal modes [9, 10, 11] of the final black hole and are dominated by the longest-lived ones. Like the black holes themselves, the quasinormal modes depend on the gravity theory. Therefore accurate ringdown measurements have the potential to put stringent bounds on alternative gravity theories.
@@
The static black holes of this theory were obtained by Sotiriou and Zhou [22, 23], and the rotating black holes by Delgado et al. [24] (see also Sullivan et al. [25]). Delgado et al. mapped carefully the domain of existence of rotating shift-symmetric EsGB black holes. This domain must, of course, be respected in the calculation of the quasinormal modes, since beyond this domain no such black holes exist. The quasinormal modes of rapidly rotating shift-symmetric EsGB black holes were obtained recently by Chung and Yunes by applying perturbation theory in the coupling constant to quadratic order in $\alpha$ [26, 27]. We here present the full results, based on an exact treatment of the dependence on the coupling constant $\alpha$ both for the background solutions and the perturbations, and then compare these exact results with the perturbative results. We note that black hole-neutron star mergers in this theory were also recently studied [28].
To obtain the quasinormal modes we employ our recently developed spectral scheme [29, 30, 31]. After testing this scheme extensively for Kerr black holes [29], we applied it to the rapidly rotating black holes of Einstein-dilaton-Gauss-Bonnet (EdGB) theory, a string theory motivated alternative theory of gravity [32, 33]. Its hairy black holes are well-studied [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], and also the emission of gravitational waves during the inspiral, merger and ringdown of binary systems was recently addressed [52].
@@ 2.2 Stationary rapidly rotating background black holes
Since the solutions possess a $Z_{2}$ symmetry $w.r.t.$ reflection on the equatorial plane, calculations can be restricted to the angular range $20\leq\theta\leq\pi/$ 2 by imposing on the equatorial plane the boundary conditions
@@
The domain of existence of black holes in shift-symmetric EsGB theory is essential to know, since beyond this domain the theory does no longer support black holes. Perturbation theory in the coupling constant $\alpha$ yields results that are not limited per se, however, they deteriorate strongly towards this boundary and become meaningless beyond the boundary. Therefore we demonstrate this domain of rotating background solutions in figure 1, which shows the dimensionless angular momentum $J/M^{2}$ versus the dimensionless coupling constant $\alpha/M^{2}$ [24].
@@ 3.2 Equations and boundary conditions
It is convenient to introduce a new set of functions, that trivially redefine the upper set of functions (see e.g. [29]) to simplify the equations and to fix the gauge. Three pairs of functions are then each expressed by a single function, reducing the number of independent functions by three. Thus we retain a set of seven independent functions which we collect in a function vector $\vec{X}$.
@@
$x=\frac{r-r_{H}}{r_{H}+1}\,,\,\,\,y=\cos\theta\,.$ (19)
The radial coordinate is thus compactified, $0\leq x\leq 1$, with the horizon located at $x=0$, and asymptotic infinity at $x=1$. The angular coordinate $y$ resides in $-1\leq y\leq 1$, with the symmetry axis located at $y=\pm 1$. In terms of the function vector $\vec{X}$ and the new coordinates $x$ and $y$ we then obtain a lengthy system of seven linear homogeneous partial differential equations, which we abbreviate as follows
@@
Next we turn to the boundary conditions, since we must ensure, that the perturbation functions are purely outgoing waves at infinity and purely ingoing waves at the horizon. To this end we factor out of the perturbation functions the common factor $\exp{i\omega R^{*}}$ with eigenvalue $\omega$ and radial function $R^{*}$, requiring the proper choice of $R^{*}$ at the boundaries of the integration. Towards radial infinity we thus require $\frac{dR^{*}}{dr}=1+\frac{2M}{r}+\mathcal{O}\left(\frac{1}{r^{2}}\right)$, and towards the horizon $\frac{dR^{*}}{dr}=\frac{g_{1}}{(r-r_{H})}+\mathcal{O}(1)$, and the $g_{1}$ parameter is obtained from the horizon expansion of the background functions. The boundary conditions can then be given in operator form, e.g., they read at infinity
$\mathcal{A}_{\mathrm{I}}(x,y)\vec{X}(x,y)|_{x=1}=0,\,\,\,\quad\mathrm{I}=1,...,7\,,$ (21)
where $\mathcal{A}_{\mathrm{I}}$ are linear operators in $(x,y)$, and analogously for the other boundaries of the domain of integration.
@@ 3.3 Spectral decomposition
$\widetilde{X}_{\mathrm{I}}(x,y)$ $=$ $\sum_{k=0}^{N_{x}-1}\,\,\sum_{l=|M_{z}|}^{N_{y}+|M_{z}|-1}C_{{\mathrm{I}},k,l}\,T_{k}(x)\,P_{l}^{M_{z}}(y)\,(1-y^{2})^{-M_{z}/2}\,,$ (22)
for ${\mathrm{I}}=1,...,7$. The parameters $N_{x}$ and $N_{y}$ specify the truncation of the expansions and the number of grid points chosen in the corresponding coordinate domains. The double expansion then introduces a set of constants $C_{{\mathrm{I}},k,l}$ for each of the seven perturbation functions.
We choose the grid by employing Gauss-Lobato points for the radial coordinate $x$ and an equidistant mesh for the angular coordinate $y$. To evaluate the perturbation functions $X_{I}$ at the grid points, we need to interpolate the background functions and their derivatives at these grid points. This leads us finally to a system of $7\times N_{x}\times N_{y}$ algebraic equations to determine the constants $C_{{\mathrm{I}},k,l}$. These algebraic equations can be put in the following form,
@@ 4 Spectrum of quasinormal modes for rotating shift-symmetric EsGB black holes
In figure 2 we show the real and imaginary parts of the modes scaled by mass, respectively, $M\omega_{R}$ and $M\omega_{I}$, as a function of the scaled coupling constant $\xi=\alpha/M^{2}$. The modes are computed for a range of scaled angular momenta $j=J/M^{2}=0$, 0.2, 0.4, 0.6, 0.8 which includes the static modes for this theory. The dark khaki vertical lines in the figures indicate the limiting scaled coupling constant for each $j$, which decreases as $j$ grows. These limiting values are derived from the domain of existence of the shift-symmetric EsGB black holes.
@@
Figure 3: Comparison of the exact modes with $l=M_{z}=2$ fundamental polar-led and axial-led quasinormal modes where a weak coupling (wc) approximation was employed [26], for $j=0.2,0.4,0.6,0.8$. Also shown is a comparison between the static modes with a full coupling and the perturbative modes for a small scaled angular momentum $j=0.005$ (topmost) [26].
@@
As the black hole rotates faster, the corresponding limiting scaled coupling constant (indicated by the vertical line in figure 2) decreases. Of all the cases we have considered, the largest deviation between our results and the wc ones is found in figure 3 in the scaled imaginary part of the axial modes for $j=0.8$ at about $\xi=0.168$, which amounts to a deviation of about 55%. This shows that the effect of the exact treatment of the coupling constant becomes highly relevant for rapidly rotating black holes. On the other hand, the scaled real part of the perturbative axial modes agrees well with our results up to about $\xi=0.071$ (i.e., 30% of the limiting scaled coupling constant). Due to a deterioration of our numerical precision, we exhibit the modes for $j=0.8$ only up to $\xi=0.168$, that is about 70% of the limiting scaled coupling constant.
@@
Figure 4: Comparison between the $l=M_{z}=2$ fundamental polar-led and axial-led modes of the shift-symmetric EsGB black holes (ss) and EdGB black holes (dil) [31], for $j=0.2,0.8$.
Lastly, let us compare the exact modes obtained here for the shift-symmetric EsGB theory with those obtained for EdGB theory in [31]. Figure 4 shows a comparison between the $l=M_{z}=2$ fundamental polar-led and axial-led modes for $j=0.2$ and $j=0.8$, respectively. The pink vertical lines with a smaller value of the scaled coupling constant denote the limiting $\xi$ values for EdGB, while the dark khaki vertical lines denote those for the shift-symmetric EsGB. We find that the modes of these two theories with a different coupling function agree well for smaller coupling constants. This can be expected since the shift-symmetric EsGB has a linear coupling function, that corresponds to the first (relevant) term in the expansion of the EdGB coupling function.

### v2

@@ 1 Introduction
There are various reasons to expect that GR will be superseded by a more complete theory of gravity. Thus in recent years numerous alternative theories of gravity have been suggested, driven largely by cosmological considerations and attempts to reconcile GR with quantum mechanics [6, 7]. Of course, any viable gravitational theory should conform with all experiments and observations [8]. But currently the strong gravity sector is still in need of far more testing [7]. Here the next generation of gravitational wave detectors should play an important role.
After the merger of two black holes the newly formed highly excited black hole undergoes ringdown where it emits damped sinusoidal gravitational waves, while it evolves towards a stationary state. These waves are associated with the quasinormal modes [11, 12, 13] of the final black hole and are dominated by the longest-lived ones (although nonlinearities in the black-hole ringdown could also be important [9, 10]). Like the black holes themselves, the quasinormal modes depend on the gravity theory. Therefore accurate ringdown measurements have the potential to put stringent bounds on alternative gravity theories (see e.g. [14, 15]).
@@
The static black holes of this theory were obtained by Sotiriou and Zhou [26, 27], and the rotating black holes by Delgado et al. [28] (see also Sullivan et al. [29]). Delgado et al. mapped carefully the domain of existence of rotating shift-symmetric EsGB black holes. This domain must, of course, be respected in the calculation of the quasinormal modes, since beyond this domain no such black holes exist. The quasinormal modes of rapidly rotating shift-symmetric EsGB black holes were obtained recently by Chung and Yunes by applying perturbation theory in the coupling constant to quadratic order in $\alpha$ [30, 31] (see also [32] for a recent study on the quasinormal modes of the static configurations). We here present the full results, based on an exact treatment of the dependence on the coupling constant $\alpha$ both for the background solutions and the perturbations, and then compare these exact results with the perturbative results. We note that black hole-neutron star mergers in this theory were also recently studied [33].
To obtain the quasinormal modes we employ our recently developed spectral scheme [34, 35, 36, 37]. After testing this scheme extensively for Kerr black holes [34], we applied it to the rapidly rotating black holes of Einstein-dilaton-Gauss-Bonnet (EdGB) theory, a string theory motivated alternative theory of gravity [38, 39]. Its hairy black holes are well-studied [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], and also the emission of gravitational waves during the inspiral, merger and ringdown of binary systems was recently addressed [58].
@@ 2.2 Stationary rapidly rotating background black holes
Since the solutions possess a $Z_{2}$ symmetry $w.r.t.$ reflection on the equatorial plane, calculations can be restricted to the angular range $0\leq\theta\leq\pi/2$ by imposing on the equatorial plane the boundary conditions
@@
It is essential to know the domain of existence of black holes in shift-symmetric EsGB theory since beyond this domain the theory no longer supports black holes. Perturbation theory in the coupling constant $\alpha$ yields results that are not limited per se, however, they deteriorate strongly towards this boundary and become meaningless beyond the boundary. Therefore we demonstrate this domain of rotating background solutions in figure 1, which shows the dimensionless angular momentum $J/M^{2}$ versus the dimensionless coupling constant $\alpha/M^{2}$ [28].
@@ 3.1 Ansatz
The background metric and scalar field are given by (6) and (7) respectively. A more thorough description of the background solutions can be found in [28].
@@ 3.2 Equations and boundary conditions
It is convenient to redefine the set of functions $\{a_{1},a_{2},a_{3},a_{4},N_{0},H_{1},L_{0},T_{0},S_{0}\}$ in (14) and (15) to simplify the equations and to fix the gauge (see e.g. [34]). Three pairs of the metric functions can be redefined using three functions, hence reducing the number of independent metric functions by three. The redefinitions are
$a_{1}(r,\theta)$ $=$ $-iM_{z}\frac{h_{0}(r,\theta)}{\sin{\theta}}\,$, (19)
$a_{2}(r,\theta)$ $=$ $\sin{\theta}\,\partial_{\theta}h_{0}(r,\theta)\,$, (20)
$a_{3}(r,\theta)$ $=$ $-iM_{z}\frac{h_{1}(r,\theta)}{\sin{\theta}}\,$, (21)
$a_{4}(r,\theta)$ $=$ $\sin{\theta}\,\partial_{\theta}h_{1}\,$, (22)
$N_{0}(r,\theta)$ $=$ $\left(g^{(bg)}_{rr}(r,\theta)\right)^{-1}N(r,\theta)\,$, (23)
$L_{0}(r,\theta)$ $=$ $\left(g^{(bg)}_{rr}(r,\theta)\right)L(r,\theta)\,$, (24)
$T_{0}(r,\theta)$ $=$ $\left(g^{(bg)}_{\theta\theta}(r,\theta)\right)T(r,\theta)\,$, (25)
$S_{0}(r,\theta)$ $=$ $\left(g^{(bg)}_{\phi\phi}(r,\theta)\right)T(r,\theta)\,$, (26)
in addition to the function $H_{1}$ that remains. Thus we retain a set of seven independent functions which we collect in a function vector $\vec{X}=[\widetilde{H}_{1},\widetilde{T},\widetilde{N},\widetilde{L},\widetilde{h}_{0},\widetilde{h}_{1},\widetilde{\Phi}_{1}]$ that is composed of six of the metric perturbation functions plus the scalar one, and where we have used the following parametrization,
$H_{1}$ $=$ $\widetilde{H}_{1}(x,y)\,\frac{1}{x(1-x)}\,e^{i\hat{R}}\,$, (27)
$T$ $=$ $\widetilde{T}(x,y)\,e^{i\hat{R}}\,$, (28)
$N$ $=$ $\widetilde{N}(x,y)\,\frac{1}{1-x}\,e^{i\hat{R}}\,$, (29)
$L$ $=$ $\widetilde{L}(x,y)\,\frac{1}{x^{2}(1-x)}\,e^{i\hat{R}}\,$, (30)
$h_{0}$ $=$ $\widetilde{h}_{0}(x,y)\,\frac{1}{1-x}\,e^{i\hat{R}}\,$, (31)
$h_{1}$ $=$ $\widetilde{h}_{1}(x,y)\,\frac{1}{x(1-x)}\,e^{i\hat{R}}\,$, (32)
$\Phi$ $=$ $\widetilde{\Phi}_{1}(x,y)\,(1-x)\,e^{i\hat{R}}\,$, (33)
where the function $\hat{R}$ is chosen such that the perturbations satisfy the ingoing/outgoing wave conditions that are explained below.
@@
$x=\frac{r-r_{H}}{{r}+1}\,,\,\,\,y=\cos\theta\,.$ (34)
The radial coordinate is thus compactified, $0\leq x\leq 1$, with the horizon located at $x=0$, and asymptotic infinity at $x=1$. The angular coordinate $y$ resides in $1\leq y\leq 1$, with the symmetry axis located at $y=\pm 1$. In terms of the function vector $\vec{X}$ and the new coordinates $x$ and $y$ we then obtain a lengthy system of seven linear homogeneous partial differential equations, which we abbreviate as follows
@@
Next we turn to the boundary conditions, since we must ensure, that the perturbation functions are purely outgoing waves at infinity and purely ingoing waves at the horizon. To this end we factor out of the perturbation functions the common factor $\exp{(i\omega R^{*})}$ with eigenvalue $\omega$ and radial function $R^{*}$, requiring the proper choice of $R^{*}$ at the boundaries of the integration, as follows. At infinity, the perturbation functions are described by
$T$ $=$ $e^{i\omega R^{*}}\left(T^{+}(\theta)+\mathcal{O}\left(\frac{1}{r}\right)\right)\,$, (36)
$H_{1}$ $=$ $re^{i\omega R^{*}}\left(H^{+}_{1}(\theta)+\mathcal{O}\left(\frac{1}{r}\right)\right)\,$, (37)
$N$ $=$ $re^{i\omega R^{*}}\left(N^{+}(\theta)+\mathcal{O}\left(\frac{1}{r}\right)\right)\,$, (38)
$L$ $=$ $re^{i\omega R^{*}}\left(L^{+}(\theta)+\mathcal{O}\left(\frac{1}{r}\right)\right)\,$, (39)
$h_{0}$ $=$ $re^{i\omega R^{*}}\left(h^{+}_{0}(\theta)+\mathcal{O}\left(\frac{1}{r}\right)\right)\,$, (40)
$h_{1}$ $=$ $re^{i\omega R^{*}}\left(h^{+}_{1}(\theta)+\mathcal{O}\left(\frac{1}{r}\right)\right)\,$, (41)
$\Phi_{1}$ $=$ $\frac{1}{r}e^{i\omega R^{*}}\left(\Phi_{1}^{+}(\theta)+\mathcal{O}\left(\frac{1}{r}\right)\right)\,$, (42)
and at the horizon,
$T$ $=$ $e^{-i(\omega-2\Omega_{H})R^{*}}\left(T^{-}(\theta)+\mathcal{O}\left(r-r_{H}\right)\right)\,$, (43)
$H_{1}$ $=$ $\frac{r_{H}}{r-r_{H}}e^{-i(\omega-2\Omega_{H})R^{*}}\left(H^{-}_{1}(\theta)+\mathcal{O}\left(r-r_{H}\right)\right)\,$, (44)
$N$ $=$ $e^{-i(\omega-2\Omega_{H})R^{*}}\left(N^{-}(\theta)+\mathcal{O}\left(r-r_{H}\right)\right)\,$, (45)
$L$ $=$ $\frac{r_{H}^{2}}{(r-r_{H})^{2}}e^{-i(\omega-2\Omega_{H})R^{*}}\left(L^{-}(\theta)+\mathcal{O}\left(r-r_{H}\right)\right)\,$, (46)
$h_{0}$ $=$ $e^{-i(\omega-2\Omega_{H})R^{*}}\left(h^{-}_{0}(\theta)+\mathcal{O}\left(r-r_{H}\right)\right)\,$, (47)
$h_{1}$ $=$ $\frac{r_{H}}{r-r_{H}}e^{-i(\omega-2\Omega_{H})R^{*}}\left(h^{-}_{1}(\theta)+\mathcal{O}\left(r-r_{H}\right)\right)\,$, (48)
$\Phi_{1}$ $=$ $\frac{1}{r}e^{-i(\omega-2\Omega_{H})R^{*}}\left(\Phi_{1}^{-}(\theta)+\mathcal{O}\left(r-r_{H}\right)\right)\,$, (49)
where $\Omega_{H}=\frac{\sqrt{M^{2}-4r_{H}^{2}}}{2M(M+2r_{H})}$. Towards radial infinity we thus require $\frac{dR^{*}}{dr}=1+\frac{2M}{r}+\mathcal{O}\left(\frac{1}{r^{2}}\right)$, and towards the horizon $\frac{dR^{*}}{dr}=\frac{g_{1}}{(r-r_{H})}+\mathcal{O}(1)$, where the $g_{1}$ parameter is obtained from the horizon expansion of the background functions and corresponds to the inverse surface gravity $\kappa_{H}^{-1}$ as in the case of Kerr black holes [60], the perturbative shift-symmetric EsGB black holes [30, 31] and EdGB black holes [36, 37], i.e. $g_{1}=\kappa_{H}^{-1}$. After introducing these expressions into the field equations, it is then straightforward to obtain the boundary conditions that the perturbation functions have to satisfy at each boundary. These boundary conditions can then be given in operator form, e.g., they read
$\mathcal{A}_{\mathrm{I}}(x,y)\vec{X}(x,y)|_{x=0}=0,\,\,\,\quad\mathrm{I}=1,...,7\,$, (50)
$\mathcal{B}_{\mathrm{I}}(x,y)\vec{X}(x,y)|_{x=1}=0,\,\,\,\quad\mathrm{I}=1,...,7\,$, (51)
where $\mathcal{A}_{\mathrm{I}},\mathcal{B}_{\mathrm{I}}$ are linear operators in $(x,y)$ (see [37]).
@@ 3.3 Spectral decomposition
$\widetilde{X}_{\mathrm{I}}(x,y)$ $=$ $\sum_{k=0}^{N_{x}-1}\,\,\sum_{l=|M_{z}|}^{N_{y}+|M_{z}|-1}C_{{\mathrm{I}},k,l}\,T_{k}(x)\,P_{l}^{M_{z}}(y)\,$, (52)
for ${\mathrm{I}}=1,...,7$. The functions $\widetilde{X}_{\mathrm{I}}=[\widetilde{H}_{1},\widetilde{T},\widetilde{N},\widetilde{L},\widetilde{h}_{0},\widetilde{h}_{1},\widetilde{\Phi}_{1}]$ as seen from (27)-(33) are those that remain after a re-parametrization of the perturbation functions where the appropriate ingoing ($\sim e^{-i(\omega-2\Omega_{H})R^{*}}$) and outgoing ($\sim e^{i\omega R^{*}}$) behaviors have been factored out, hence ensuring that our resulting perturbation system has the correct wave behavior at the horizon and at infinity. The parameters $N_{x}$ and $N_{y}$ specify the truncation of the expansions and the number of grid points chosen in the corresponding coordinate domains. The double expansion then introduces a set of constants $C_{{\mathrm{I}},k,l}$ for each of the seven perturbation functions, where $k$ comes from the expansion in $x$ and $l$ from the expansion in $y$.
We choose the grid by employing Gauss-Lobato points for the radial coordinate $x$
$x_{I}=\frac{1}{2}\left(1+\cos\left(\frac{I-1}{N_{x}-1}\pi\right)\right)\,,\quad I=1,...,N_{x}\,$, (53)
and an equidistant mesh for the angular coordinate $y$
$y_{K}=2\left(\frac{K-1}{N_{y}-1}\right)-1\,,\quad K=1,...,N_{y}\,.$ (54)
To evaluate the perturbation functions $X_{I}$ at the grid points, we need to interpolate the background functions and their derivatives at these grid points. In addition, from the 11 field equations for the perturbation functions, we only need to choose 7, while the remaining 4 can be used to test the accuracy of the solutions. This leads us finally to a system of $7\times N_{x}\times N_{y}$ algebraic equations to determine the constants $C_{{\mathrm{I}},k,l}$. These algebraic equations can be put in the following form,
@@
There are several ways to estimate the accuracy of the quasinormal modes calculated. For example, since from the 11 field equations only 7 need to be used to obtain the solutions, the remaining 4 are used to estimate the accuracy. In addition, other auxiliary parameters of the spectral decomposition can be varied to estimate the error (for example, the original number of points of the numerical background, the collocation grid, etc). We make this analysis for each of the modes calculated, and for all the results reported in this work. The relative precision of the quasinormal modes is always better than $10^{-5}$. See [34, 36, 37, 62] for extensive details of the method.
@@ 4 Spectrum of quasinormal modes for rotating shift-symmetric EsGB black holes
In figure 2 we show the real and imaginary parts of the modes scaled by mass, respectively, $M\omega_{R}$ and $M\omega_{I}$, as a function of the scaled coupling constant $\xi=\alpha/M^{2}$. The modes are computed for a range of scaled angular momenta $j=J/M^{2}=0$, 0.2, 0.4, 0.6, 0.8 which includes the static modes for this theory. The dark khaki vertical lines in the figures indicate the limiting scaled coupling constant for each $j$, which decreases as $j$ grows. These limiting values are derived from the domain of existence of the shift-symmetric EsGB black holes. In the Appendix we provide the corresponding fitting functions for the modes in tables 2 - 11.
@@
Figure 3: Comparison of the exact modes with $l=M_{z}=2$ fundamental polar-led and axial-led quasinormal modes where a weak coupling (WC) approximation was employed [30], for ${J/M^{2}}=0.2,0.4,0.6,0.8$. Also shown is a comparison between the static modes with a full coupling and the perturbative modes for a small scaled angular momentum ${J/M^{2}}=0.005$ (topmost) [30].
@@
As the black hole rotates faster, the corresponding limiting scaled coupling constant (indicated by the vertical line in figure 2) decreases. Of all the cases we have considered, the largest deviation between our results and the WC ones is found in figure 3 in the scaled imaginary part of the axial modes for $j=0.8$ at about $\xi=0.168$, which amounts to a deviation of about 55%. This shows that the effect of the exact treatment of the coupling constant becomes highly relevant for rapidly rotating black holes. On the other hand, the scaled real part of the perturbative axial modes agrees well with our results up to about $\xi=0.071$ (i.e., 30% of the limiting scaled coupling constant). We summarize in table 1 the threshold of the coupling constant beyond which the perturbative computations [30, 31] depart from our exact results. Due to a deterioration of our numerical precision, we exhibit the modes for $j=0.8$ only up to $\xi=0.168$, that is about 70% of the limiting scaled coupling constant. The decline in the precision of the modes is mainly due to the decrease in the numerical precision of the background solutions, that deteriorates as we get closer and closer to the limit value of the coupling constant. In particular, for this case, the estimated mean error of the background solutions is less than $10^{-5}$. More precisely, it is around $10^{-8}$ close to GR, and around $10^{-6}$ close to the limiting coupling constant. Discussions about the limiting behavior of the solutions can be found in [28]. As we discussed above, the precision of all the modes obtained in this work is of $10^{-5}$ or better.
Threshold of $\alpha/M^{2}$
Axial Polar
$J/M^{2}$ Real freq. Im. freq. Real freq. Im. freq.
$\leq$ 0.005 0.23 0.15 0.12 0.10
0.2 0.27 0.25 0.11 0.25
0.4 0.19 0.12 0.11 0.15
0.6 0.24 0.12 0.09 0.09
0.8 0.02 0.05 0.03 0.01
Table 1: Threshold of the scaled coupling constant $\alpha/M^{2}$ for each scaled angular momentum $J/M^{2}$ for the scaled real frequency $M\omega_{R}$ and imaginary frequency $M\omega_{I}$ of the $l=M_{z}=2$ fundamental axial-led and polar-led quasinormal modes, derived from the mode comparison in figure 3. Beyond the threshold, results from the weak coupling approximation begin to deviate from exact results. Thresholds in the first row are obtained from a comparison between our exact static modes and the perturbative modes for $J/M^{2}=0.005$.
@@
Figure 4: Comparison between the $l=M_{z}=2$ fundamental polar-led and axial-led modes of the shift-symmetric EsGB black holes ($s$) and EdGB black holes (dil) [36, 37], for ${J/M^{2}}=0.2,0.8$.
Lastly, let us compare the exact modes obtained here for the shift-symmetric EsGB theory with those obtained for EdGB theory in [36, 37]. Figure 4 shows a comparison between the $l=M_{z}=2$ fundamental polar-led and axial-led modes for $j=0.2$ and $j=0.8$, respectively. The pink vertical lines with a smaller value of the scaled coupling constant denote the limiting $\xi$ values for EdGB, while the dark khaki vertical lines denote those for the shift-symmetric EsGB. We find that the modes of these two theories with a different coupling function agree well for smaller coupling constants. This can be expected since the shift-symmetric EsGB has a linear coupling function, that corresponds to the first (relevant) term in the expansion of the EdGB coupling function.
In this work, we have examined only a fraction of the modes that we typically obtain using our method. By fixing $M_{z}$, the spectral method generally captures a spectrum of modes composed of higher $l$’s ($l\geq M_{z}$) as well as higher excitation modes. For instance, for the first excitation modes of the same $l$, their real parts are usually close to those of the fundamental modes, while their imaginary parts can be of an order of $10^{-1}$ larger in absolute value compared to the fundamental modes.